OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = (x – 4)(2x – 1)(x + 7)?

OpenStudy (inkyvoyd):

x-4=0 2x-1=0 x+7=0

OpenStudy (anonymous):

(zeros are where the graph crosses the x-axis) So because each part in brackets is multiplied by another part, if any are zero f(x) is zero and therefore the graph is zero at that x value.